OpenStudy (anonymous):

A die has one red face, two blue faces, and three green faces. It is rolled 5 times. Find the chance that the red face appears on one of the rolls and the remaining rolls are green.

OpenStudy (anonymous):

This is how I calculated it: \[5\times1/6\times3/6^{4}=0,15625\] Could someone double check it (and correct me if I am wrong)? Thx.

OpenStudy (anonymous):

@Nurali, thanks for posting final result, but could you please provide the whole calculation? I would like to identify the mistake I made. Thx

OpenStudy (gorv):

(1/6*3/6*3/6*3/6*3/6)*5=5*81/6^5 5 is multiplied bcoz 1/6 can com at any postion and their are 5 positions

OpenStudy (anonymous):

Thx gorv, I solved it the same way you did, but I obviously made a mistake when I was multiplying numerators of the fractions, which led to the wrong result.

OpenStudy (anonymous):

Hi njux! So... as formula was above, you can see calculation result: 0.052 Good luck!